# Segment as You Wish: Free-Form Language-Based Segmentation for Medical Images

### v1

@@ 3.2 Free-Form Language Segmentation for Medical Images
For free-form anatomy-informed text prompts, the text encoder must learn embeddings that group similar organ segmentation intents together while clearly separating unrelated intents in distinct semantic clusters. We adopt the CLIP [44] as the foundation of text encoder for its capability of understanding semantics. Given a text prompt $p\in\mathcal{P}_{x}$ associated with the image $x$, the CLIP text encoder converts it into an embedding vector $\mathbf{t}_{p}$ in a shared embedding space: $\mathbf{t}_{p}=\text{Encoder}^{\mathcal{P}}(p)\in\mathbb{R}^{D}$, where $D$ is the dimensionality of the text embedding space. To further strengthen the model’s ability to differentiate between organ segmentation, we introduce an intention head on top of the text embeddings by CLIP. This head is a linear layer $\mathbf{W}_{\text{cls}}\in\mathbb{R}^{C\times D}$, where $C=24$ is the number of organ class. The intention logits $\mathbf{y}_{p}$ are derived for each encoded vector $\mathbf{t}_{p}$: $\mathbf{y}_{p}=\mathbf{W}_{\text{cls}}\mathbf{t}_{p}+\mathbf{b}_{\text{cls}}$. Given a corresponding medical image embedding $z^{x}$, we train the model by following loss function:
$L=\operatorname*{arg\,min}_{\{\mathbf{W}_{\text{cls}},\mathbf{b}_{\text{cls}},\mathbf{W}^{E},\mathbf{W}^{D},\mathbf{W}^{P}\}}\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\frac{1}{|\mathcal{P}_{x}|}\sum_{p\in\mathcal{P}_{x}}[\mathcal{L}_{\text{Dice}}(\hat{m}_{x}^{p},m_{x}^{p})+\mathcal{L}_{\text{ce}}(\hat{m}_{x}^{p},m_{x}^{p})+\mathcal{L}_{\text{ce}}(\mathbf{y}_{p},l_{p})]$ (1)
where $\hat{m}_{x}^{p}=\text{Decoder}(z_{x},\mathbf{t}_{p})$ and $m_{x}^{p}$ are predicted and ground truth masks. $l_{p}\in[0,...,23]$ is the ground truth organ class for the prompt. $\mathbf{W}^{E}$, $\mathbf{W}^{D}$ and $\mathbf{W}^{P}$ represent the image encoder, decoder and CLIP text encoder weights, respectively. We use both Dice loss $\mathcal{L}_{\text{Dice}}$ and cross-entropy loss $\mathcal{L}_{\text{ce}}$ for predicted masks. The classification loss $\mathcal{L}_{\text{ce}}(\mathbf{y}_{p},l_{p})$ encourages the model to correctly classify organs based on text prompts, ensuring the text embedding aligns with the intended organ class.
@@
3.3 Semantics-aware Canonicalization Learning
Figure 4: The architecture of our proposed model FLanS. First, given a set of free-form text prompts $p_{1}...,p_{n}$, the text encoder gets the text embedding, and then passes through the learned Intention Head Layer that maps the embedding to a space with explicit intention probabilities, which is useful for the FLanS model weight updating as in Eq. 1. Second, we have trained a Canonicalized Network that transforms any medical image with arbitrary orientation into a canonicalization space, making sure the encoded image aligns with the standard clinical practice to avoid ambiguity. Third, the encoded prompts (either spatial info such as Point, Bbox, or Free-form text data), together with the encoded image, will be processed with mask decoder and output the expected masks.
We incorporate roto-reflection symmetry [11] into our architecture for two key reasons: 1) Organs and anatomical structures can appear in various orientations and positions due to differences in patient positioning, imaging techniques, or inherent anatomical variations. Equivariance ensures that the model’s segmentation adapts predictably to transformations of the input image. 2) We aim to ensure our model reliably interprets and segments organs that have positional terms in their names, such as “left” or “right kidney” from text prompts regardless of the scan’s orientation, thereby enhancing the model’s robustness and accuracy.
Following [25, 41], we train a separate canonicalization network ${h}:\mathcal{X}\mapsto G$, where $\mathcal{X}$ represents the medical image sample space, $G$ represents the desired group, and $h$ is equivariant to $G$. This network generates group elements that transform input images into canonical frames, standardizing the image orientation before applying the prediction function. The Eq. 2 shows how this canonicalization process maps the transformed input back to a common space where the segmentation prediction network $p$ operates,
$f(x)=\rho_{\text{out}}({h}(x))\;p(\rho_{\text{in}}({h}^{-1}(x))x,\;\textbf{t})$ (2)
Where $p$ is the segmentation prediction network (composed of the Image Encoder and Mask Decoder in Figure 11), t is the text prompt embedding produced by our text encoder, and $\rho_{\text{in}}$ and $\rho_{\text{out}}$ are input and output representations. The segmented images or masks produced by $p$ can be transformed back with $\rho_{\text{out}}({h}(x))$ as needed. Without this transformation, $f$ is invariant; otherwise, it is equivariant. Thus, the FLanS architecture visualized in Figure 11 is invariant. We use ESCNN [5] to build the canonicalization network. This approach has the advantage of removing the constraint from the main prediction network and placing it on the network that learns the canonicalization function. Appendix A provides a detailed introduction of symmetry and equivariant networks.
As the entire architecture achieves invariance or equivariance through canonicalization, the model produces the same segmentation or consistently transforms the segmentation according to the transformed input. In other words, the model always segment the same areas of interest regardless of the image’s orientation with the same text prompt. For example, as long as the ground truth “right kidney” mask of a CT image has been shown to the model once, no matter how the orientation of the CT image and the location of the right kidney changes, the model will always segment the same area.
However, without proper training, ${h}(x)$ might map different images to inconsistent canonical frames, causing a distribution shift in the inputs to the prediction network and affecting performance. Thus, training the canonicalization network togther with the segmentation prediction network is essential to ensure consistent mapping to the desired frame. It is worth noting that users can choose to disable the canonicalizer when working with anatomy-agnostic prompts, as the segmented organ may differ if the original image is not in the canonical frame. The decision depends on whether the user wants to segment the original or the canonicalized image, as the model will segment whatever image is fed into the image encoder based on the provided text prompts.
3.4 Training Strategy
We employ a three-stage training strategy for FLanS: 1) Learning canonicalization: we train the canonicalization network independently using FLARE22 training samples applied with random transformations from the O(2) group. The network is optimized using MSE loss between the canonicalized samples and their original counterparts. This encourages the canonicalization network to map transformed samples back to their canonical orientations as seen in the FLARE22 dataset, preventing it from selecting arbitrary orientations that could degrade the performance of the prediction network. 2) Learning text-prompted segmentation: we train FLanS with the queries from Generator $\mathbf{G}$ as introduced in Section 3.1, without the canonicalization network on the original scans, using both anatomy-informed and anatomy-agnostic prompts. This ensures that the segmentation network learns to respond accurately to different types of prompts without interference from canonicalization and data augmentation. 3) Learning augmentation and alignment: In the final stage, we perform joint training on all scans, applied with random O(2) transformations. Since the canonicalization network may not always generate the exact canonical orientation the segmentation network is accustomed to in the beginning, this serves as a form of free augmentation for the segmentation networks. Over time, the canonicalization and segmentation networks align.
4 Experiment
4.1 Datasets and Experiments Setup
Image Datasets
To develop an effective organ segmentation model, we collected 1,437 CT scans from 7 public datasets, covering 24 partially labeled organs. Of these, 1,089 scans from MSD [1], BTCV [19], WORD [37], AbdomenCT-1K [40], FLARE22 [39], and CHAOS [26] are used for training. The rest 65 scans, consisting of 10% of the FLARE22 dataset (in-domain), the official validation set of WORD (in-domain), and the official test set of RAOS [36] (out-of-domain), were used to evaluate model performance. To standardize the quality and reduce domain gaps across datasets, we applied pre-processing techniques such as slice filtering and intensity scaling to all CT scans. The finalized dataset comprised 91,344 images for training and validation, and 9,873 for testing. Detailed information on the dataset statistics and pre-processing steps are in Appendix B.
Text Datasets
Our text dataset was constructed using two types of queries: anatomy-agnostic and anatomy-informed. First, for each image, we identified organs corresponding to 6 representative positions: leftmost, rightmost, topmost, bottom, smallest, and largest. For each of these 6 position indicators, 100 anatomy-agnostic queries were generated, resulting in a set of 600 queries to serve as anatomy-agnostic segmentation prompts. [^3] To ensure accurate position-to-organ mapping, position-driven organ-agnostic queries were applied only to images containing more than nine labeled organs during training. Second, for each organ, we generated 480 anatomy-informed queries in an expertise-driven style using the RAG query generator. By combining both anatomy-agnostic and anatomy-informed queries, we formed a text dataset comprising 12,120 unique queries for model training. During testing, a comprehensive text set was used, containing both in-domain and out-of-domain queries. Specifically, we generated 30 RAG-generated expertise-style queries (25%, in-domain), 30 human-generated non-expertise-style queries (25%, out-of-domain), and 60 RAG-generated non-expertise-style queries (50%, out-of-domain) for each organ, forming a test set of 120 queries per organ and 2,880 queries across all organs. Detailed information on the generation of the text queries is in Appendix B.
Experiment Setup
All experiments were conducted on an AWS ml.p3dn.24xlarge instance equipped with 8 V100 GPUs, each with 32 GB of memory. We used a batch size of 16 and applied the CosineAnnealingLR learning rate scheduler, initializing the learning rate for all modules at 0.0001. The AdamW optimizer was employed for training. A small $D_{8}$ equivariant canonicalization network was used, consisting of 3 layers, a hidden dimension of 8, and a kernel size of 9. To maintain consistency across the input and output formats, all scans from different datasets were resized to 1024$\times$1024 and both predicted and ground truth masks were resized to 256$\times$256 for fair comparison. For images with a single channel, the channel was duplicated to 3. All models’ performance on the test sets is reported using both the Dice coefficient [51] and normalized surface distance [22] .
Figure 5: Left: Segmentation with anatomy-informed prompts. We could observe that FLanS can precisely segment the organ described in free-form text prompts, while other baselines make mistakes in identifying the organs. Right: Segmentation with anatomy-agnostic prompts. We could observe that the FLanS is texture-aware, descriptions of the sizes and positions can be understood, and is competitively accurate to the direct Bbox segment.
4.2 Anatomy-informed Segmentation
@@
Table 2: Anatomy-Informed Segmentation Results: FLanS consistently outperforms baselines on both organ name and free-form text prompts segmentation tasks, demonstrating superior language understanding and segmentation accuracy across in-domain and out-of-domain datasets, even when applied with random transformations.
@@
As for the baselines, the Universal Model [33] is the only published medical imaging foundation model that considers free-form text descriptions. This model integrates text description embeddings during training, while segmentation at the testing and inference stages is performed using organ IDs. Consequently, we evaluate this model with prompts consisting solely of organ names. Another widely used approach for text-prompt segmentation involves combining CLIP-based models [44] with segmentation models [32, 54]. In these methods, segmentation models first generate potential masks based on a set of random bounding box or point prompts that span the entire image. CLIP-based models then embed both the text prompt and the cropped images from these masks. The final mask is selected based on the highest similarity between the cropped image embedding and the text embedding. To cover this approach, we include two additional baselines: 1) CLIP + MedSAM, where MedSAM [62] is SAM [28] fine-tuned on medical imaging datasets; and 2) MedCLIP + MedSAM, where MedCLIP [59], a contrastive learning framework trained on diverse medical image-text datasets, is paired with MedSAM for segmentation.
@@
4.3 Anatomy-Agnostic Segmentation
Table 3: Anatomy-Agnostic Segmentation Results: Comparison of FLanS using positional and size information text prompts vs. MedSAM and SAM2 using Bboxes or points. FLanS achieves competitive or superior performance across both in-domain and out-of-domain test sets.
@@ 4.3 Anatomy-Agnostic Segmentation
SAM2-large (Point-prompt) 0.585 0.652 0.534 0.551 0.488 0.497
SAM2-large (Bbox-prompt) 0.873 0.906 0.848 0.802 0.818 0.749
MedSAM (Bbox-prompt) 0.887 0.872 0.783 0.781 0.697 0.681
FLanS (Free-form text) 0.844 0.841 0.855 0.853 0.851 0.850
To evaluate our model’s ability to understand anatomy-agnostic text prompts, we tested its segmentation performance using prompts that contain only positional or size-related information. To the best of our knowledge, no existing model is designed to handle anatomy-agnostic text prompts. Therefore, we chose state-of-the-art MedSAM [62] (SAM fine-tuned on medical imaging datasets) and the latest SAM2 [46] as baselines. However, instead of text prompts, these models were provided with ground-truth organ Bboxes or point prompts. Our goal in this experiment is for FLanS to achieve comparable results to the baselines because FLanS is only given text prompts with positional or size information while the baselines are given the bounding box or point prompts of ground truth organ.
As shown in Table 3, FLanS the best or second-best performance across both in-domain and out-of-domain test sets. MedSAM performs well on the FLARE and WORD test sets but struggles on the RAOS test set due to the lack of training on that dataset. SAM2, when provided with bounding box prompts, consistently performs well across all test sets and demonstrates strong generalizability. However, its performance significantly degrades with point prompts, likely because medical scans lack the distinct edges present in the datasets SAM2 was originally trained on. The right panel of Fig. 5 visualizes the segmentations produced by the best baseline and FLanS, along with their corresponding anatomy-agnostic text prompts. It demonstrates that FLanS can reliably segment the correct organs based on the provided positional or size information, such as largest and lower right.
4.4 Ablation Study on the Model Architecture
Figure 6: The model without canonicalization incorrectly highlights the left kidney due to confusion between anatomical position (“right kidney”) and the organ’s appearance on the right side of the image.
Table 4: Ablation study: prediction performance of FLanS and its variants with progressively removed components on the FLARE22 original and transformed test sets. Each row represents a version of the model with one additional component removed.

### v2

@@ 3.2. Free-Form Language Segmentation
For anatomy informed text prompts, the text encoder must learn embeddings that group similar organ segmentation intents together while clearly separating unrelated intents in distinct semantic clusters. We adopt the CLIP (Radford et al., 2021) as the foundation of text encoder for its capability of understanding semantics. Given a text prompt $p\in\mathcal{P}_{x}$ associated with the image $x$, the CLIP text encoder converts it into an embedding vector $\mathbf{t}_{p}$ in a shared embedding space: $\mathbf{t}_{p}=\text{Encoder}^{\mathcal{P}}(p)\in\mathbb{R}^{D}$, where $D$ is the dimensionality of the text embedding space.
To further strengthen the model’s ability to differentiate between organ segmentation, we introduce an intention head on top of the text embeddings by CLIP. This head is a linear layer $\mathbf{W}_{\text{cls}}\in\mathbb{R}^{C\times D}$, where $C=24$ is the number of organ class. The intention logits $\mathbf{y}_{p}$ are derived for each encoded vector $\mathbf{t}_{p}$: $\mathbf{y}_{p}=\mathbf{W}_{\text{cls}}\mathbf{t}_{p}+\mathbf{b}_{\text{cls}}$. Given a corresponding medical image embedding $z^{x}$, we train the model by following loss function:
(1) $L=\mathop{\arg\min}_{\mathcal{W}^{*}}\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\frac{1}{|\mathcal{P}_{x}|}\sum_{p\in\mathcal{P}_{x}}\big{[}\mathcal{L}_{\text{Dice}}(\hat{m}_{x}^{p},m_{x}^{p})+\mathcal{L}_{\text{ce}^{*}}\big{]}$
where $\mathcal{W}^{*}$ are sets of model weights: $\mathcal{W}^{*}=\{\mathbf{W}_{\text{cls}},\mathbf{b}_{\text{cls}},\mathbf{W}^{E},\mathbf{W}^{D},\mathbf{W}^{P}\}$, and $\hat{m}_{x}^{p}=\text{Decoder}(z_{x},\mathbf{t}_{p})$ and $m_{x}^{p}$ are predicted and ground truth masks. $l_{p}\in[0,...,23]$ is the ground truth organ class for the prompt. $\mathbf{W}^{E}$, $\mathbf{W}^{D}$ and $\mathbf{W}^{P}$ represent the image encoder, decoder and CLIP text encoder weights, respectively. We use both Dice loss $\mathcal{L}_{\text{Dice}}$ and cross-entropy loss $\mathcal{L}_{\text{ce}^{*}}$ for predicted masks. And for $\mathcal{L}_{\text{ce}^{*}}$:
(2) $\mathcal{L}_{\text{ce}^{*}}=\mathcal{L}_{\text{ce}}(\hat{m}_{x}^{p},m_{x}^{p})+\mathcal{L}_{\text{ce}}(\mathbf{y}_{p},l_{p}).$
it contains two parts, apart from measuring the predicted mask $\hat{m}_{x}^{p}$ compared to the ground truth mask $m_{x}^{p}$, the second term $\mathcal{L}_{\text{ce}}(\mathbf{y}_{p},l_{p})$ encourages the model to correctly classify organs based on text prompts, ensuring the text embedding aligns with the intended organ class.
@@
Table 2. Anatomy-Informed Segmentation Results: FLanS consistently outperforms baselines on both organ name and free-form text prompts segmentation tasks, demonstrating superior language understanding and segmentation accuracy across in-domain and out-of-domain datasets, even when applied with random transformations.
@@
3.3. Semantics-Aware Canonicalization Learning
We incorporate roto-reflection symmetry (Cohen and Welling, 2016) into our architecture for two key reasons: 1) Organs and anatomical structures can appear in various orientations and positions due to differences in patient positioning, imaging techniques, or inherent anatomical variations. Equivariance ensures that the model’s segmentation adapts predictably to transformations of the input image. 2) We aim to ensure our model reliably interprets and segments organs that have positional terms in their names, such as “left” or “right kidney” from text prompts, regardless of the scan’s orientation, thereby enhancing the model’s robustness and accuracy.
Following (Kaba et al., 2022; Mondal et al., 2023), we train a separate canonicalization network ${h}:\mathcal{X}\mapsto G$, where $\mathcal{X}$ represents the medical image sample space, $G$ represents the desired group, and $h$ is equivariant to $G$. This network generates group elements that transform input images into canonical frames, standardizing the image orientation before applying the prediction function.
(3) $f(x)=\rho_{\text{out}}({h}(x))\;p(\rho_{\text{in}}({h}^{-1}(x))x,\;\textbf{t})$
The Eq. 3 shows how this canonicalization process maps the transformed input back to a common space where the segmentation prediction network $p$ operates, where $p$ is the segmentation prediction network (composed of the Image Encoder and Mask Decoder in Figure 13), t is the text prompt embedding produced by our text encoder, and $\rho_{\text{in}}$ and $\rho_{\text{out}}$ are input and output representations. The segmented images or masks produced by $p$ can be transformed back with $\rho_{\text{out}}({h}(x))$ as needed. Without this transformation, $f$ is invariant; otherwise, it is equivariant. Thus, the FLanS architecture visualized in Figure 13 is invariant. We use ESCNN (Cesa et al., 2022) to build the canonicalization network. This approach has the advantage of removing the constraint from the main prediction network and placing it on the network that learns the canonicalization function. Appendix A provides a detailed introduction of symmetry and equivariant networks.
As the entire architecture achieves invariance or equivariance through canonicalization, the model produces the same segmentation or consistently transforms the segmentation according to the transformed input. In other words, the model always segments the same areas of interest regardless of the image’s orientation with the same text prompt. For example, as long as the ground truth “right kidney” mask of a CT image has been shown to the model once, no matter how the orientation of the CT image and the location of the right kidney change, the model will always segment the same area.
However, without proper training, ${h}(x)$ might map different images to inconsistent canonical frames, causing a distribution shift in the inputs to the prediction network and affecting performance. Thus, training the canonicalization network together with the segmentation prediction network is essential to ensure consistent mapping to the desired frame. It is worth noting that users can choose to disable the canonicalizer when working with anatomy-agnostic prompts, as the segmented organ may differ if the original image is not in the canonical frame. The decision depends on whether the user wants to segment the original or the canonicalized image, as the model will segment whatever image is fed into the image encoder based on the provided text prompts.
3.4. Training Strategy
We employ a three-stage training strategy for FLanS: 1) Learning canonicalization: The canonicalization network is trained independently using FLARE22 training samples with random transformations from the O(2) group. It is optimized with MSE loss between the canonicalized samples and their original counterparts, encouraging the network to map transformed samples back to their canonical orientations in the FLARE22 dataset and avoid arbitrary orientations that could impair prediction performance. 2) Learning text-prompted segmentation: We train FLanS with queries from Generator $\mathbf{G}$ as introduced in Section 3.1, using both anatomy-informed and anatomy-agnostic prompts on original scans, ensuring the segmentation network accurately responds to different prompts without interference from canonicalization. 3) Learning augmentation and alignment: In the final stage, we perform joint training on all scans with random O(2) transformations. As the canonicalization network may not always provide the exact canonical orientation expected by the segmentation network initially, this serves as free augmentation, gradually aligning the canonicalization and segmentation networks.
4. Experiment
4.1. Datasets and Experiments Setup
Image Datasets: We collected CT scans covering 24 labeled organs from the datasets below. Of these, 1,089 scans from MSD (Antonelli et al., 2022), BTCV (Gibson et al., 2018), WORD (Luo et al., 2021), AbdomenCT-1K (Ma et al., 2022), FLARE22 (Ma et al., 2023), and CHAOS (Kavur et al., 2019) were used for training, while 65 scans from FLARE22, WORD, and the RAOS (Luo et al., 2024) test set were used for evaluation. Pre-processing steps (slice filtering, intensity scaling) standardized quality and reduced domain gaps, resulting in 91,344 images for training/validation and 9,873 for testing. Details on dataset statistics and pre-processing steps are provided in the Appendix B. Text Datasets: Our text dataset includes 12,120 training queries (600 anatomy-agnostic, 11,520 anatomy-informed) generated by the RAG query generator. For testing, we created 2,880 queries (120 per organ) covering both in-domain and out-of-domain queries. Experiment Setups: All experiments were conducted on AWS instance equipped with 8 V100 GPUs, each with 32 GB of memory, more details in Appendix B and  C.
Figure 6. Left: Segmentation with anatomy-informed prompts. We could observe that FLanS can precisely segment the organ described in free-form text prompts, while other baselines make mistakes in identifying the organs. Right: Segmentation with anatomy-agnostic prompts. We could observe that the FLanS is texture-aware, descriptions of the sizes and positions can be understood, and is competitively accurate to the direct Bbox segment.
Figure 7. Left: t-SNE visualization of the free-form text prompt embedding. FLanS (Ours) can effectively distinguish between different organ-related queries. Right: Canonicalized CT scans from $D_{4}$ and $D_{8}$ canonicalization networks for a batch of randomly transformed scans from the FLARE22 dataset. Medical images can be successfully transformed back to an aligned canonicalization space.
4.2. Anatomy-Informed Segmentation
@@
Among existing baselines, BiomedParse (Zhao et al., 2024) is, to our best knowledge, the only model designed to handle free-form text prompts. While Universal Model (Liu et al., 2023b) is also a medical imaging foundation model that incorporates text descriptions during training, it performs segmentation at the inference stage using organ IDs. Consequently, we evaluate this model with prompts consisting solely of organ names. Another widely used approach for text-prompt segmentation involves combining CLIP-based models (Radford et al., 2021) with segmentation models (Li et al., 2024; Wang et al., 2024). In these methods, segmentation models first generate potential masks based on a set of random bounding box or point prompts that span the entire image. CLIP-based models then embed both the text prompt and the cropped images from these masks. The final mask is selected based on the highest similarity between the cropped image embedding and the text embedding. To cover this approach, we include two additional baselines: 1) CLIP + MedSAM, where MedSAM (Wu et al., 2023) is SAM (Kirillov et al., 2023) fine-tuned on medical imaging datasets; and 2) MedCLIP + MedSAM, where MedCLIP (Wang et al., 2022d), a contrastive learning framework trained on diverse medical image-text datasets, is paired with MedSAM for segmentation. We include a SOTA few-shot learning model, UniverSeg (Butoi* et al., 2023), to ensure our baselines remain competitive. Unlike the other methods, UniverSeg does not support text prompts. Instead, we evaluate it using a few support samples with their corresponding ground truth organ masks, rather than prompts based on organ names.
@@
4.3. Anatomy-Agnostic Segmentation
Table 3. Anatomy Agnostic Segmentation Results: Comparison of FLanS using positional and size information text prompts vs. MedSAM and SAM2 using Bboxes or points. FLanS achieves competitive or superior performance across both in-domain and out-of-domain test sets.
@@ 4.3. Anatomy-Agnostic Segmentation
SAM2-large (Point) 0.585 0.652 0.534 0.551 0.488 0.497
SAM2-large (Bbox) 0.873 0.906 0.848 0.802 0.818 0.749
MedSAM (Bbox) 0.887 0.872 0.783 0.781 0.697 0.681
FLanS (Free-form) 0.844 0.841 0.855 0.853 0.851 0.850
To evaluate our model’s ability to understand anatomy-agnostic text prompts, we tested its segmentation performance using prompts that contain solely positional or size-related information. To the best of our knowledge, no existing model is designed to handle anatomy-agnostic text prompts. Therefore, we chose state-of-the-art MedSAM (Wu et al., 2023) (SAM fine-tuned on medical imaging datasets) and the latest SAM2 (Ravi et al., 2024) as baselines. However, instead of text prompts, these models were provided with ground-truth organ Bboxes or point prompts. Our goal in this experiment is for FLanS to achieve comparable results to the baselines because FLanS is only given text prompts with positional or size information while the baselines are given the bounding box or point prompts of ground truth organ.
As shown in Table 3, achieves top performance across in-domain and out-of-domain test sets. MedSAM performs well on the FLARE and WORD test sets but struggles on the RAOS due to the lack of training on that dataset. SAM2, when provided with bounding box prompts, consistently performs well across all test sets and shows strong generalizability. However, its performance degrades with point prompts, likely because medical scans lack the distinct edges present in the datasets SAM2 was originally trained on. The right panel of Fig. 6 visualizes the segmentations produced by the best baseline and FLanS, along with anatomy-agnostic text prompts. FLanS reliably segments the correct organs based on the provided positional or size descriptions.
4.4. Ablation Study on the Model Architecture
Table 4. Ablation study: prediction performance of FLanS and its variants with progressively removed components on the FLARE22 original and transformed test sets. Each row represents a version of the model with one additional component removed.